An airplane flying at a rate of 500 mph leaves Hawaii en route to Alaska. One-half hour later, a second airplane leaves Hawaii in the same direction flying at a rate of 600 miles per hour. How long will it take the second airplane to overtake the first?
Answer by checkley77(12844)

D=RT
D=500T
D=600(T-1)
500T=600(T-1)
500T=600T-600
500T-500T=-600
-100T=-600
T=-600/-100
T=6 HOURS AFTER THE FIRST PLANE LEAVES THEY WILL MEET.
PROOF:
500*6=600(6-1)
3,000=600*5
3,000=3,000